OpenStudy (anonymous):

solve the equation by factoring: 2x-17=9/x

OpenStudy (anonymous):

Solve the whole thing with x: x(2x-17=9/x) 2x^2 -17x = 9 ----> 2x^2 - 17x - 9 = 0, then factor it!

OpenStudy (anonymous):

(2x-9)(x+1)

OpenStudy (anonymous):

(2x+1)(2x-9)

OpenStudy (anonymous):

do you know the quadratic formula?

OpenStudy (anonymous):

i dont think its possible to do factoring with this one

OpenStudy (anonymous):

oops sorry! its (2x+1)(x-9)